OpenStudy (anonymous):

Is a triangle with sides of length 6 ft, 21 ft, 23 ft a right triangle? Explain.

OpenStudy (apoorvk):

check if the sum of the square of the two least sides equals the square of the longest side --> basically check if this satisfies the pythagoras theorem.

OpenStudy (amorfide):

a²+b²=c² 6²+21²=36 + 441 = 477 square root 477=21.8 so this is not a right angle triangle as a²+b² is not equal to c²